OpenStudy (help_people):

A firecracker shoots up from a hill 150 feet high, with an initial speed of 110 feet per second. Using the formula H(t) = −16t2 + vt + s, approximately how long will it take the firecracker to hit the ground? Eight seconds Nine seconds 10 seconds 11 seconds

OpenStudy (help_people):

@LiveLaughDie @Nnesha @pooja195

OpenStudy (sohailiftikhar):

what is s?

OpenStudy (help_people):

im not sure @sohailiftikhar

OpenStudy (sohailiftikhar):

it should be some constant value like 5,3 etc u checked the statement clearly ?

OpenStudy (sohailiftikhar):

take is 5 ..now out height H=0 to find the time 0=-16t^2+vt+5

OpenStudy (sohailiftikhar):

now 16t^2-vt=5 put value of v from given data v=110fet/sec 16t^2-110t=5 divide whole equation with 16 to get simplest form

OpenStudy (help_people):

16t^2=5+110 16T^2=115 is tis correct?

OpenStudy (sohailiftikhar):

no where is the t with 110 ?

OpenStudy (help_people):

? could you just show me for my understudying @sohailiftikhar

OpenStudy (sohailiftikhar):

just find the value of t from the equation which i gave you ignor negative value and get +ve value that will be the answer

OpenStudy (help_people):

t=2.680951

OpenStudy (help_people):

@sohailiftikhar

OpenStudy (sohailiftikhar):

did u do your work correctly ? what two values you have got ? for t

OpenStudy (help_people):

-2.680951

OpenStudy (help_people):

there ya goooo @sohailiftikhar

OpenStudy (sohailiftikhar):

ok it's correct

OpenStudy (help_people):

ok I'm still not sue what the answer is :/ @sohailiftikhar

OpenStudy (sohailiftikhar):

lol we are not sure about s which we taken as 5 ..so take it easy it's 2 sec

OpenStudy (help_people):

that not one of the answers :( @sohailiftikhar

OpenStudy (sohailiftikhar):

so what is answer ?O.o

OpenStudy (help_people):

10? @sohailiftikhar

OpenStudy (sohailiftikhar):

lol what ? no it can be possible it only can 2 or near it your answer is correct i'm sure ..from where u get this huh ?

OpenStudy (help_people):

here are the options: Eight seconds Nine seconds 10 seconds 11 seconds

Nnesha (nnesha):

setup the equation when it hits the ground h(t)=0 n v=110 and s=150 now plug in the numbers and solve the qudratic equation

OpenStudy (help_people):

ok

OpenStudy (help_people):

so im solving for s?

Nnesha (nnesha):

no solve for t they already gave you v and s value

OpenStudy (help_people):

what are they ?

Nnesha (nnesha):

\(\color{blue}{\text{Originally Posted by}}\) @Nnesha setup the equation when it hits the ground h(t)=0 n v=110 and s=150 now plug in the numbers and solve the qudratic equation \(\color{blue}{\text{End of Quote}}\) read this

OpenStudy (help_people):

ook so what would the formula be to solve for t ?

Nnesha (nnesha):

\[\huge\rm \frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\] quadratic formula

Nnesha (nnesha):

a , b ,c values are \[\huge\rm Ax^2+Bx+C=0\] A= leading coefficient b=middle term c= constant term